Minutes of the 8th Committee Meeting of the Dublin University Rifle Club - Freshers Week Meeting

Moderator Steven Watterson Minutes taker Mark Dennehy Present Steven Watterson (Captain),

Benjamin Jones (Treasurer),

Mark Dennehy (Secretary/Match Secretary), John Keeney (Ammo&Equipment Officer), Alan Smyth (Chief Range Officer)

Absent(with apologies) Conor McDermottroe (Public Relations Officer)

Absent Jason Sibley, Deirdre Ninaber (Ladies’ Coordinator)

Date 10 September 2001 1800 – 2030 Distribution DURC Committee

Overview of topics

1 email 1

2 Freshers’ Week Preperations 1

1 email

A discussion on the rifle.club@tcd.ie address was had. John is to forward email to Steven from this address despite the high level of spam in the email. Mark doesn’t think he is getting the email from John’s redirect.

2 Freshers’ Week Preperations

It was decided to start shooting as soon as possible in term, beginning the first week of term. Cathy’s Demo week is concurrent with this and so it was discussed. Ben is unsure what DUCAC is planning as regards this and so will talk to Cathy about it.

2 Freshers’ Week Preperations

Task: Find out about the demo week from Cathy (Ben) At this point

Ben noted that “We’re Cool”.

The venue for the demo week was thought to be problematic as it was out-of-the-way and the timing clashed with the start of lectures.

John suggested we push for membership from second-year and higher students. Postgrads for example, are not usually in clubs. Steven pointed out that that was due to postgrads having lives and committments allready. Ben suggested we get the GSU to remind postgrads about the clubs available to them instead of a hard sell approach, as those avoiding Freshers Week are doing so for a reason. Steven pointed out that covering other gates is a drain on resources as we need four people on the stand.

We also need to find out the days for registration for different faculties to try to predict manpower requirements for busy days (ie. Science/Engineering days).

The Safety Meeting was discussed. It was agreed to have it as a basic introductory meeting, to keep it short (under 20 minutes) and to have it on time as last year saw a theatre full of people and no club people until Ben arrived late. It was decided to bill the meeting as an Introductory Meeting to make it seem more friendly.

The range needs more decoration as it seems a bit too focused at present and not congenial to relaxation. This was decided to be something that could be discussed later however.

The Introductory Meeting was to be scheduled for as soon as possible - Friday of Freshers’ Week was thought to be too conflicting a date as most people have social engagements, so Monday thereafter was selected. Mark is to book the theatre. Task: Book a lecture hall for the Introductory Meeting (Mark)

The content was to be Basic Safety, a demonstration of standing and prone shooting, which Steven volunteered for (Mark having done so last year), we would show targets and photographs and so forth, and then retire to the range for the reception.

The reception’s expenses came up as an important point. Beer for this function is expensive - on the order of 10-12 crates were consumed last year. Selling the beer was discussed to offset costs, but Mark pointed out that DUCAC would frown on this so the idea was ruled out.

Membership fees were discussed. Steven was concerned that we not price ourselves out of the market. Ben says that we have to find out what the DUCAC levy is this year and change according to that. Task: Find out what the DUCAC levy will be this year (Ben)

Steven suggested different fees for new and returning members, with returning members paying a higher fee - this was however opposed on the grounds that the return would be minimal due to our turnover, but the higher fee would not be taken well by casual shooters. The idea was dropped when the return turned out to be on the order of £20.

Mark noted that we need to be lobbying DUCAC for a higher grant now with the data collected on expenses for the past year. This was agreed with, but no specific tasks were identified.

Regular social meetings were discussed as being a sorely needed feature in the Club’s setup. Steven suggested that during Freshers’ Week that we meet every night in a pub and have an open invitation for members to join us there. Despite plees of financial and liver-based concern, this was agreed as necessary. It was decided to announce this with a sign on the stand when Mark pointed out that it did sound like a lame chat-up line to mention it when lobbying for members. It was decided that this would then be a regular social event during term. Logistical discussions ensued - the final decision was that a set time (Thursdays at 2000) would be announced, and that off-duty ROs and members would meet there, to be joined by the on-duty ROs as time permitted. The on-duty ROs would remind members shooting that night that they could retire to the Pavillion bar afterwards for this event.

Membership cards were discussed and agreed on as a good idea, though without the photographs from last year as the camera was an extra security risk on the stand. Alan is to try printing on card from his office printers to try to reduce the cost of printing the cards. Task: Experiment with printing membership cards (Alan)

The stand rota and the range rota for the week were to be drawn up by the CRO. All ROs are expected to pull duty during the week. Task: Draw up range and stand rotas (Alan)

Steven proposed the idea of opening the range during Freshers’ week for introductory shooting. This was agreed to be an excellent idea, and after some debate it was agreed to shoot Air Rifle, the pellet cost being included in the membership fee as it was negligible. The idea was that someone would join at the stand, get their membership card and come down to the range, where they would get to shoot air rifle under supervision (one RO to a shooter, and the RO loads the rifle). Only two people to shoot at a time to allow the ROs to monitor on a one-to-one basis. The range was to open as the stand opened and close as it closed (after the stand closed, there could be a problem with students drinking and wanting to shoot afterwards).

Squad training was deferred from discussion as Steven has been thinking about reorganising the squads and wants more time to arrange his thoughts.

The topic of sporter rifle was raised - should we keep it for the coming year? Alan was for keeping it as it was fun, easy, and therefore good for beginners. Ben agreed as people liked the discipline. Mark said he had no problem with supporting it but only if the Club was serious and attended competitions - that otherwise it was a waste of time and resources. After spirited debate it was agreed that we would keep sporter, giving it a shared day with smallbore rifle, and that we would go on one training day and one competition as a trial during the year. It was however, stated that smallbore shooting was to begin with prone rifle to learn to shoot and then progression to standing shooting could be made for reasons of safety. RO training in sporter was also required.

Competitions were then discussed, specifically the procedures for inviting people to attend. Limiting attendance to those we could carry by car on a squad-based system was discussed and dropped. It was agreed that limited teams would attend competitions in international venues such as EARC, Comber, XB, or the Isle of Man, but ’local’ competitions at RRPC, FSC, MRC, DRC, BRC, UCDRC and WTSC were all reachable by bus and so were not a problem for extra people. However, a basic competency level is to be required before attending, and a Code of Conduct will have to be signed by each attendee.

The scheduling of disciplines was discussed and the decision was as follows: Monday Air Rifle Tuesday Prone Rifle Wednesday Air Rifle Thursday Prone Rifle/Sporter Rifle Friday Seasonal : Air Rifle during Air Rifle season Prone Rifle during Prone season

Ammunition needs to be ordered. A long discussion was held on this topic and that of financial arrangements. The problem is that we need to raise nearly £800 in seed money to purchase 10,000 rounds of T22 and Master ammunition - thereafter sales of that ammunition can raise the money to buy the next batch. A loan from DUCAC was ruled out as it would need to be a recurring loan and not a one-time arrangement, and it was not possible to be sure that the loan would be paid according to the academic year timetable. No method of raising that capital was thought of though. Mark suggested we ask for sponsorship from ammunition companies in the form of ammunition, though this was a short-term solution. Task: Talk to Millards and other ammunition dealers regarding prices (John) Task: Send John a copy of his list of dealers (Steven)

Three-position Air Rifle was brought up. WTSC and the PDF are planning on introducing this discipline and UCDRC have been considering it. It is a useful 3P training tool and Steven and Mark support it’s introduction but it will have to wait until later in the year due to the extra RO training required.

Steven brought up the .22 blocks. He felt they were not necessary. This was disagreed with on the grounds that it does give a weeks grace to a rank beginner to acclimitize to the rifle, trigger, sights and the general feel of shooting. Also, sling adjustment for a pain-free position is a nontrivial task and not best attempted from the beginning. However, it was agreed a replacement was needed - Mark pointed out that the current blocks are a safety hazard, both from hernias and ricochets. It was agreed a simple rest was a better solution, and possible designs were discussed but nothing was agreed upon.

Financial concerns were revisited. Steven compiled a list of the expenses incurred in our social events and expressed concern at the total. It was agree that costs must be cut

- functions should run on a BYOB footing for smaller functions such as the grudge match with the DU fencing club. The Committee lunch was however, defended as relatively sacrosant - however Roly’s Bistro was ruled out as a venue ☺. 2020: John,

Alan leave

RO Credits and finances were discussed. Ben agreed to include RO Credit reports in his weekly Treasurers report. Penalties and limits on credit were discussed but no final decision made other than the decision that a change is badly needed due to abuse of the RO Credit system by the ROs. Suggested remedies included charging interest on credits above a threshold or refusing more credit than a set level. The problem of enforcing these rules was discussed.

T.B.N.O.B.T.M.W.D.C.

2 Freshers’ Week Preperations Steven Watterson, Club Captain Mark Dennehy, Club Secretary

Tasks agreed on at meeting

Ben Find out about the demo week from Cathy (10 September 2001) [/143] /2 Mark Book a lecture hall for the Introductory Meeting (10 September

2001) [/182] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Ben Find out what the DUCAC levy will be this year (10 September

2001) [/200] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Alan Experiment with printing membership cards (10 September 2001) [/237] /2 Alan Draw up range and stand rotas (10 September 2001) [/243] . . . . . /2 John Talk to Millards and other ammunition dealers regarding prices (10

September 2001) [/317] . . . . . . . . . . . . . . . . . . . . . . . . . . . /2

Steven Send John a copy of his list of dealers (10 September 2001) [/318] . /2